Points P, E,and A are marked on a circle whose center is R. In quadrilateral PEAR, angles A and E are found to be 54 degrees and 113 degrees, respectively. What are the other two angles?
Answer by rothauserc(4718)

measure of angle E = (1/2) of arc PA
:
113 = (1/2) of arc PA
:
arc PA = 226 degrees
:
therefore
:
measure of angle PRA = 360 - 226 = 134 degrees
:
measure of angle RPE = 360 - 113 - 134 - 54 = 59 degrees
:
***************************************************
The other two angles are 134 degrees and 59 degrees
***************************************************
: